OpenStudy (anonymous):

PLEASE HELP ME IDONTUNDERSTAND LOL solve :1+(1/x^2)=(3/x)

OpenStudy (anonymous):

1+(1/x^2)=(3/x) x(1+(1/x^2))=(3/x)x

OpenStudy (anonymous):

would xx^2 work as well?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

x(1+(1/x^2))=(3/x)x distribute x+1/x=3

OpenStudy (anonymous):

wait..i dont get it.. isn't this a quad. formula?

OpenStudy (anonymous):

oh, i get the problem...

OpenStudy (anonymous):

1-3/x+1/x^2=0

OpenStudy (anonymous):

so..if i use the quad. formula my answer would be... -3(+or-)rad5 over 2?

OpenStudy (anonymous):

i'm assuming yes, my hands are full of food.

OpenStudy (anonymous):

...cause i redid the problem and i got 3(+or-)rad5 over 2

OpenStudy (anonymous):

im not sure if its -3 or 3 :(

OpenStudy (anonymous):

i get 2.618

OpenStudy (anonymous):

then its 3 now need to find my mistake.. thank you very much!! :)

OpenStudy (anonymous):

very awkward question.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

show me what you did. maybe i can help?

OpenStudy (anonymous):

i made a reallly really stupid mistake..i thought the b (the 1st b) in the quad. formula was just b....its a negative b...

OpenStudy (anonymous):

a hundred face palms would not be enough